# Characterizing GPU Resilience and Impact on AI/HPC Systems

### v3

@@ 3.2. Data Processing Pipeline
Input : Error logs with timestamps $E=\{(e_{1},t_{1}),\dots,(e_{n},t_{n})\}$, regex patterns =$\mathcal{R}=\{r_{1},r_{2},\dots\}$, time window $\Delta t$
@@
$E^{\prime}\leftarrow\emptyset$ // Initialize output set
@@
      $E_{r}\leftarrow\{(e_{i},t_{i})\in E\mid e_{i}\text{ matches }r\}$ // Filter errors with regex
       $i\leftarrow 1$ // Loop through errors in a matched group
@@
             // Loop through later errors within the matched group
             while $i+1\leq|E_{r}|$ do
@@
                   // Error has identical message and is close in time
                   if $e_{\text{next}}=e_{\text{first}}$ and $t_{\text{next}}-t_{\text{latest}}\leq\Delta t$ then
                         $t_{\text{latest}}\leftarrow t_{\text{next}}$ // Discard latest error
                        $i\leftarrow i+1$
@@
Error Statistic Metrics. Using the coalesced error logs as input, the pipeline computes standard error statistics metrics such as the count and the MTBE as in (Oles et al., 2024) (Figure 4, item (2)). For errors that are not directly countable from XID logs, we estimate their occurrence by correlating related XIDs within the same time interval (Oles et al., 2024; NVIDIA, 2024; NVIDIA Corporation, 2024). In particular, the number of uncorrectable ECC memory errors can be inferred by summing up the number of RREs and RRFs, as they are mutually exclusive outcomes of an uncorrectable ECC memory error recovery event. Subsequently, the number of consecutive SBEs can be obtained by subtracting the number of DBEs from the number of uncorrectable ECC memory errors.
In addition, we computed system-wide MTBE and derived per-node MTBE by normalizing the error count using the number of GPU nodes in Delta. The per-node MTBE indicates the operational hours a single Delta GPU node can function before encountering an error. For GPU memory resilience characterization, we additionally derived per-GPU MTBE by normalizing the error count using the number of GPU in a node and per-GB MTBE by normalizing the error count with per GPU memory capacity in GB. The per-GPU MTBE reflects the operational hours a single GPU can function before encountering an error.
Error Propagation Analysis. We performed error propagation analysis (Figure 4, item (3)) to capture how errors propagate within a GPU and across different GPUs while measuring the propagation time. The propagation probability from GPU error $e_{1}$ to $e_{2}$ is defined:
$P(e_{2}|e_{1})=\frac{\#e_{2}}{\text{Total }\#e_{1}},\quad t_{e_{2}}-t_{e_{1}}\leq\Delta t.$
A propagation path is created if $e_{2}$ occurs immediately after $e_{1}$ within a predefined time window $\Delta t$. If there is no succeeding error $e_{2}$ after $e_{1}$ within $\Delta t$, then $e_{1}$ is a terminal error that does not propagate. For intra-GPU [^7] GPU devices are identified by their node ID and PCI Express bus address. propagation, we require errors $e_{1}$ and $e_{2}$ to be on the same GPU device, whereas for the inter-GPU propagation, $e_{1}$ and $e_{2}$ are from two distinct GPUs on the same node. We recorded the time difference between the initial ($e_{1}$) and subsequent errors ($e_{2}$), referred to as the propagation time, for each propagation event. A shorter propagation time suggests a higher correlation between $e_{1}$ and $e_{2}$. We applied the same $\Delta t$ selection criteria as the error coalescing analysis and selected the $\Delta t=5$ seconds in error propagation analysis.
@@
  • *Row remapping, contained memory error, and uncontained memory error are new resilience features introduced starting with the NVIDIA Ampere architecture for uncorrectable ECC memory error management.
  • *Per-node MTBE in hours is derived by multiplying system MTBE by the number of GPU nodes of the corresponding GPU type. The number of nodes and number of GPUs per node type are specified in Figure 2.
@@ 4. Characterizing GPU Resilience
This section characterizes and compares the resilience of Delta’s NVIDIA A100 (Ampere) and H100 (Hopper) GPUs. Specifically, we discuss error statistics and error propagation of GPU errors in three categories: (i) GPU memory, (ii) GPU hardware, and (iii) NVLink interconnect, as described in Section 2.2 and Table 1. These errors are critical because they interrupt user jobs and lead to unplanned node downtime, as we show in Section 5. We first highlight error statistics and key findings from our analysis and then focus on error propagation for each of the three error categories on Delta’s workload. Note that we do not directly compare Delta with Blue Waters (Di Martino et al., 2014), Titan (Nie et al., 2016), or Summit (Oles et al., 2024) because of the significant generational gap from the GPUs used in those machines.
@@ 4.1. Error Statistics and Result Highlights
GPU Memory. Recall that, each A100 GPU incorporates 40 GB of HBM2e memory, and each H100 GPU incorporates 96 GB of HBM3 memory. H100’s shows 3.2×3.2$\times$3.2 × lower per GPU MTBE for uncorrectable ECC memory errors compared to A100’s, despite comparable per-GB MTBE. This reduced resilience likely stems from H100’s higher memory density, which trades resilience for capacity and performance. The observed lower per node MTBE of RREs and higher RRF counts on H100 GPUs further suggest that current resilience features are insufficient for increased memory capacity. As a uncorrectable memory error can cause a multi-GPU job to fail, enabling uninterrupted execution in such cases at scale remains an open challenge.
  1. (i)
    Delta’s H100 GPUs exhibit lower memory resilience than the A100 GPUs. Specifically, the per GPU MTBE is 3.2$\times$ lower on H100 GPUs at 88,768 hours versus 283,271 hours for A100 GPUs. Based on the per GPU MTBE, we calculated the per GB MTBE to be 8,521,728 hours [^8] We calculated the per GB MTBE by multiplying the per node MTBE with the total memory of all GPUs in GB for that node. For example, in a 4-way H100 GPU node with 96 GB of memory, the per GB MTBE = $22,192\times 96\times 4=8,521,728$ hours. for H100’s HBM3 memory vs. 11,330,826 hours for A100’s HBM2e memory, a 24% reduction.
@@
    We attribute the decrease in resilience is primarily due to the higher memory capacity (96 GB vs. 40 GB, a 2.4×2.4$\times$2.4 × increase), which increases the chances of bit flips. We additionally hypothesize that H100 memory resilience is worse due to (a) a lower signaling voltage that increases susceptibility to bit flips (Wu et al., 2024) and (b) an increased number of stacks that make heat dissipation challenging and degrade the resilience of memory modules, of the HBM3 memory.
  2. (ii)
@@
    Uncorrectable ECC memory error-recovery mechanisms (e.g., row remapping and error containment, see Section 2.1) improve GPU memory resilience and reduce service interruptions (NVIDIA Corporation, 2024). The error recovery mechanisms mitigate uncorrectable memory errors with a probability of 0.92 on H100 according to error propagation analysis in Section 4.2. However, these mechanisms may not scale well with the increase in GPU memory capacity in H100. For example, the available spare rows for row remapping are capped at the same 512 rows (NVIDIA Corporation, 2024), which is not proportional to the 2.4×2.4$\times$2.4 × increase in memory capacity. Such insufficiency can be evident from the significantly lower per node MTBE of RRE on the H100 GPUs than the A100 GPUs, indicating more frequent recovery events. In addition, we observed 8 RRFs on H100 GPUs during the early operational period, which indicates memory recovery failure due to exhaustion of spared memory rows. We have yet to observe an RRF on the A100 GPUs during the much longer operational period.
@@
GPU Hardware and NVLink Interconnect. GPU hardware such as GSP, PMU SPI, and NVLink are critical components from a resilience perspective for A100 GPUs, leading to node downtime and job failures. In this context, H100 GPUs demonstrate significant improvements in GPU hardware resilience, especially in critical components such as the GSP, PMU SPI, and NVLink. We believe that these improvements are likely due to the tightly integrated heterogeneous CPU-GPU architecture of GH200 (NVIDIA Corporation, 2023) and driver-level enhancements (NVIDIA, 2023). We present a detail analysis below.
  1. (i)
@@
    Among the GPU hardware components on A100 GPUs, GSP, intended as a performance enhancement, is the most vulnerable due to its lack of robust detection and recovery. Our error propagation analysis (Section 4.2) shows that over 99% of GSP errors put the GPU in an error state and lead to job failure if encountered. A GSP error requires a GPU reset or a node reboot to recover, introducing significant overheads. In addition, PMU SPI communication errors propagate downstream and cause MMU errors with a probability of 0.88, leading to MMU errors, which then, in turn, result in a job failure over 90% of the time. Although PMU SPI communication errors are high-impact errors, they are not highlighted in NVIDIA’s developer’s manual (NVIDIA, 2024).
  2. (ii)
@@
    Despite error detection mechanisms such as CRC and recovery mechanisms such as message retransmitting, NVLink GPU interconnect errors are frequent (1,922 total NVLink errors) on A100 GPUs, with a node MTBE of 1,185 hours (system-wide MTBE of 11 hours); 42% (801) of the NVLink errors affected two or more GPUs. Although NVIDIA (NVIDIA, 2024) indicates that a GPU reset or a node reboot is required to clear NVLink, Delta SREs suggested that NVLink errors were largely benign and did not always lead to job failures. Indeed, the user job impact analysis shows that an NVLink error has a 54% chance of leading to a job failure (see Section 5). When an NVLink error is observed but does not affect a user job, it may be because NVLink is primarily used for communication rather than computation in many jobs or because multiple NVLink errors within the same job have a consolidated impact, resulting in high occurrence rates but minimal application disruption.
  3. (iii)
@@
    H100 GPUs have substantially improved GPU hardware resilience over A100 GPUs: only 3 GSP errors were observed during the measurement period. Moreover, the number of GPU Fallen off the bus errors, NVLink errors [^9] We tested the NVLink to ensure that the NVLinks were enabled., and PMU SPI errors were not observed in H100 GPUs. We attribute this improvement to (a) the better packaging and tight of CPU and GPU into a single heterogeneous compute module, which significantly reduces integration errors and enhances the resilience of the CPU-GPU complex, and (b) the NVIDIA driver upgrades appear to have improved GSP and NVLink stability (also observed by the Delta SREs).
@@
Figure 5. Intra-GPU uncorrectable memory error recovery paths in H100 GPUs. Numbers on the edges show propagation probability. The precise sub-second timing information is not available in the H100 nodes’ system logs. All propagation time for H100 GPUs memory errors were within one second.
[…]
Graph showing memory error propagation probabilities.
@@ 4.2. GPU Error Propagation
Successful Error Recovery and Containment. As shown in  Figure 5, row remapping recovery (RRE) triggered by an uncorrectable ECC memory error has a success rate of 0.59 on H100 GPUs. For the 33% of the row remapping events that fail (RRF), the GPU still contains the uncorrectable memory error because only the affected user jobs were terminated, and only the faulty page was offlined. Overall, considering both uncorrectable memory error recovery paths (RRE and error containment after an RRF), the impact of uncorrectable memory errors was alleviated 92% of the time on H100, while the GPU can remained operable. Such uninterrupted operations were not achievable on previous-generation GPUs (e.g., Kepler in (Di Martino et al., 2014; Tiwari et al., 2015a) and Volta in (Oles et al., 2024)), as an uncorrectable memory error would immediately cause user job interruption and put GPU in an error state, necessitating a GPU reset to recover (NVIDIA, 2024).
@@
Figure 6. Intra-GPU hardware error propagation probabilities in A100 GPUs. Numbers on the edges show propagation probabilities and average propagation time in seconds.
[…]
Plot showing hardware error propagation with probabilities
@@
GSP errors can be caused by either GSP firmware bugs (NVIDIA, 2023) or demanding workload. For example, Delta SREs observed that these errors were highly correlated with heavy ML benchmarks, and they suggested that GSP errors are high-impact errors whose recovery requires manual node draining and reboots. Our propagation analysis confirm that the GSP is a single point of failure on both A100 GPUs in part because of their spontaneous nature and high downstream impact (e.g., GPU hangs) on the GPU.
@@
GPU Fallen Off the Bus. GPU Fallen Off the Bus errors were logged when the GPU driver could not reach the GPU over the system bus. This error is an integration error often caused by a loose GPU-motherboard connection or contact failure because of thermal cycles (Tiwari et al., 2015a). Over 99% of the errors of this type lead to similar errors in close successions and eventually put the GPU into an error state.
Our GPU hardware error propagation analysis suggests that the error detection and recovery of GSP, PMU, and the communication interfaces (e.g., PMU SPI) need to be improved via duplications and error-detection and correction mechanisms to prevent single points of failure, as evident in A100 GPUs. In fact, AWS recommends disabling GSP for stability over performance benefits (AWS Official, 2023). By improving GSP hardware and driver software combined with a tightly integrating CPU-GPU architecture, the H100 GPUs in the GH200 Superchips significantly improve GPU hardware resilience over the A100 GPUs. Notably, apart from three GSP errors, which follow the same propagation paths as GSP errors in A100 GPUs, we observed no other hardware error propagation paths on H100 GPUs.
@@
Figure 7. NVLink intra-GPU and inter-GPU error propagation in A100 GPUs. Numbers on the edges show propagation probabilities and average propagation time in seconds.
[…]
Graph showing NVLink error propagation probabilities.
@@
Figure 8. Incidents in which GPU errors led to job failure.
[…]
Code snippet showing GPU error incidents leading to job failure.

### v4

@@ 3.2. Data Processing Pipeline
Input : Error logs with timestamps $E=\{(e_{1},t_{1}),\dots,(e_{n},t_{n})\}$, regex patterns $\mathcal{R}=\{r_{1},r_{2},\dots\}$, time window $\Delta t$
@@
$E^{\prime}\leftarrow\varnothing$ // Initialize output set
@@
 $E_{r}\leftarrow\{(e_{i},t_{i})\in E\mid e_{i}\text{ matches }r\}$ Filter errors with regex $i\leftarrow 1$
 // Loop through errors in a matched group
@@
     // Loop through later errors within the matched group while $i+1\leq|E_{r}|$ do
@@
        // Error has identical message and is close in time if $e_{\text{next}}=e_{\text{first}}$ and $t_{\text{next}}-t_{\text{latest}}\leq\Delta t$ then
           $t_{\text{latest}}\leftarrow t_{\text{next}}$ //Discard latest error $i\leftarrow i+1$
@@
Error Statistic Metrics. Using the coalesced error logs as input, the pipeline computes standard error statistics metrics including the error count and the MTBE as in (oles2024understanding) (Figure 4, item (2)). MTBE allows fair comparison between GPU types with different total operational hours. For errors that are not directly countable from XID logs, we estimate their occurrence by correlating related XIDs within the same time interval (oles2024understanding; nv_xid_doc; nv_mem_err). In particular, the number of uncorrectable ECC memory errors can be inferred by summing up the number of RREs and RRFs, as they are mutually exclusive outcomes of an uncorrectable ECC memory error recovery event. Subsequently, the number of consecutive SBEs can be obtained by subtracting the number of DBEs from the number of uncorrectable ECC memory errors.
We additionally computed system-wide MTBE and derived per-node MTBE by normalizing the error count using the number of GPU nodes in Delta. The per-node MTBE indicates the operational hours a single Delta GPU node can function before encountering an error. For GPU memory resilience characterization, we additionally derived per-GPU MTBE by normalizing the error count using the number of GPU in a node and per-GB MTBE by normalizing the error count with per GPU memory capacity in GB. The per-GPU MTBE reflects the operational hours a single GPU can function before encountering an error.
Error Propagation Analysis. We performed error propagation analysis (Figure 4, item (3)) to capture how errors propagate within a GPU and across different GPUs while measuring the propagation time. The propagation probability from GPU error $e_{1}$ to $e_{2}$ is defined as $P(e_{2}|e_{1})=\frac{\#e_{2}}{\text{Total }\#e_{1}},\quad t_{e_{2}}-t_{e_{1}}\leq\Delta t.$ A propagation path is created if $e_{2}$ occurs immediately after $e_{1}$ within a predefined time window $\Delta t$. If there is no succeeding error $e_{2}$ after $e_{1}$ within $\Delta t$, then $e_{1}$ is a terminal error that does not propagate. For intra-GPU [^7] GPU devices are identified by their node ID and PCI Express bus address. propagation, we require errors $e_{1}$ and $e_{2}$ to be on the same GPU device, whereas for the inter-GPU propagation, $e_{1}$ and $e_{2}$ are from two distinct GPUs on the same node. We recorded the time difference between the initial ($e_{1}$) and subsequent errors ($e_{2}$), referred to as the propagation time, for each propagation event. A shorter propagation time suggests a higher correlation between $e_{1}$ and $e_{2}$. We applied the same $\Delta t$ selection criteria as the error coalescing analysis and selected the $\Delta t=5$ seconds in error propagation analysis.
@@
*Row remapping, contained memory error, and uncontained memory error are new resilience features introduced starting with the NVIDIA Ampere for managing uncorrectable memory errors.
*Per-node MTBE (hrs) is derived by multiplying system MTBE by the number of GPU nodes of that GPU type. The number of nodes and GPUs per node type are specified in Figure 2.
@@ 4. Characterizing GPU Resilience
This section characterizes and compares the resilience of Delta’s NVIDIA A100 (Ampere) and H100 (Hopper) GPUs. Specifically, we discuss error statistics and error propagation of GPU errors in three categories: (i) GPU memory, (ii) GPU hardware, and (iii) NVLink interconnect, as described in Section 2.3 and Table 1. These errors are critical because they interrupt user jobs and lead to unplanned node downtime, as we show in Section 5. We first highlight error statistics and key findings from our analysis and then focus on error propagation for each of the three error categories on Delta’s workload. Note that we do not directly compare Delta with Blue Waters(di2014lessons), Titan(nie2016large), or Summit (oles2024understanding) as those systems used older GPUs lacking latest performance and resilience features central to our study, e.g., GSP, dynamic page offlining, row remapping, memory error containment, and NVLink recovery.
@@ 4.1. Error Statistics and Result Highlights
GPU Memory. Recall that, each A100 GPU incorporates 40 GB of HBM2e memory, and each H100 GPU incorporates 96 GB of HBM3 memory. H100’s shows $3.2\times$ lower per GPU MTBE for uncorrectable ECC memory errors compared to A100’s, despite comparable per-GB MTBE. This reduced resilience likely stems from H100’s higher memory density, which trades resilience for capacity and performance. The observed lower per node MTBE of RREs and higher RRF counts on H100 GPUs further suggest that current resilience features are insufficient for increased memory capacity. As a uncorrectable memory error can cause a multi-GPU job to fail, enabling uninterrupted execution in such cases at scale remains an open challenge. We present the detailed analysis below.
[…]
@@
We attribute the decrease in resilience is primarily due to the higher memory capacity (96 GB vs. 40 GB, a $2.4\times$ increase), which increases the chances of bit flips. We additionally hypothesize that H100 memory resilience is worse due to (a) a lower signaling voltage that increases susceptibility to bit flips (wu_atc_hbm) and (b) an increased number of stacks that make heat dissipation challenging and degrade the resilience of memory modules, of the HBM3 memory.
@@
(ii) Uncorrectable ECC memory error-recovery mechanisms (e.g., row remapping and error containment, see Section 2.3) improve GPU memory resilience and reduce service interruptions (nv_mem_err). The error recovery mechanisms mitigate uncorrectable memory errors with a probability of 0.92 on H100 according to error propagation analysis in Section 4.2. However, these mechanisms may not scale well with the increase in GPU memory capacity in H100. For example, the available spare rows for row remapping are capped at the same 512 rows (nv_mem_err), which is not proportional to the $2.4\times$ increase in memory capacity. Such insufficiency can be evident from the significantly lower per node MTBE of RRE on the H100 GPUs than the A100 GPUs, indicating more frequent recovery events. In addition, we observed 8 RRFs on H100 GPUs during the early operational period, which indicates memory recovery failure due to exhaustion of spared memory rows. We have yet to observe an RRF on the A100 GPUs during the much longer operational period.
@@
GPU Hardware and NVLink Interconnect. GPU hardware such as GSP, PMU SPI, and NVLink are critical components from a resilience perspective for A100 GPUs, leading to node downtime and job failures. In this context, H100 GPUs demonstrate significant improvements in GPU hardware resilience, especially in critical components such as the GSP, PMU SPI, and NVLink. We believe that these improvements are likely due to the tightly integrated heterogeneous CPU-GPU architecture of GH200 (nvidia_grace_hopper) and driver-level enhancements (nv_driver_release_notes). We present a detail analysis below.
@@
(i) Among the GPU hardware components on A100 GPUs, GSP, intended as a performance enhancement, is the most vulnerable due to its lack of robust detection and recovery. Our error propagation analysis (Section 4.2) shows that over 99% of GSP errors put the GPU in an error state and lead to job failure if encountered. A GSP error requires a GPU reset or a node reboot to recover, introducing significant overheads. In addition, PMU SPI communication errors propagate downstream and cause MMU errors with a probability of 0.88, leading to MMU errors, which then, in turn, result in a job failure over 90% of the time. Although PMU SPI communication errors are high-impact errors, they are not highlighted in NVIDIA’s developer’s manual (nv_xid_doc).
@@
(ii) Despite error detection mechanisms such as CRC and recovery mechanisms such as message retransmitting, NVLink GPU interconnect errors are frequent (1,922 total NVLink errors) on A100 GPUs, with a node MTBE of 1,185 hours (system-wide MTBE of 11 hours); 42% (801) of the NVLink errors affected two or more GPUs. Although NVIDIA (nv_xid_doc) indicates that a GPU reset or a node reboot is required to clear NVLink, Delta SREs suggested that NVLink errors were largely benign and did not always lead to job failures. Indeed, the user job impact analysis shows that an NVLink error has a 54% chance of leading to a job failure (see Section 5). When an NVLink error is observed but does not affect a user job, it may be because NVLink is primarily used for communication rather than computation in many jobs or because multiple NVLink errors within the same job have a consolidated impact, resulting in high occurrence rates but minimal application disruption.
@@
(iii) H100 GPUs have substantially improved GPU hardware resilience over A100 GPUs: only 3 GSP errors were observed during the measurement period. Moreover, the number of GPU Fallen off the bus errors, NVLink errors [^9] We tested the NVLink to ensure that the NVLinks were enabled., and PMU SPI errors were not observed in H100 GPUs. We attribute this improvement to (a) the better packaging and tight of CPU and GPU into a single heterogeneous compute module, which significantly reduces integration errors and enhances the resilience of the CPU-GPU complex, and (b) the NVIDIA driver upgrades appear to have improved GSP and NVLink stability (also observed by the Delta SREs).
GPU Errors Temporal Analysis. Temporal analysis revealed that GPU error rates followed the classic “bathtub” reliability curve. In the pre-operational period, i.e., the “infant-mortality” phase, Delta underwent extensive testing, hardware replacements, and software fixes, during which the system-wide MTBE was 0.15 hours [^10] Evaluated using the critical-errors listed in Table 1.. In the operational period, the system-wide MTBE rose $10\times$ to 1.4 hours, marking a transition to the “normal-life phase” with lowered error rate. Furthermore, estimation of the hazard-rate ($H(t)=\int h(t)\,dt$) and cumulative hazard-rate using Nelson-Aalen estimator showed that GPU hazard-rate varies over time but exhibits no clear temporal trends during the operational period (e.g., hazard rate with mean 0.7, std 7.94 for A100 GPUs).
@@ 4.2. GPU Error Propagation
Figure 5. Intra-GPU uncorrectable memory error recovery paths in H100 GPUs. Numbers on the edges show propagation probability. The precise sub-second timing information is not available in the H100 nodes’ system logs. All propagation time for H100 GPUs memory errors were within one second.:
@@
Successful Error Recovery and Containment. As shown in  Figure 5, row remapping recovery (RRE) triggered by an uncorrectable ECC memory error has a success rate of 0.59 on H100 GPUs. For the 33% of the row remapping events that fail (RRF), the GPU still contains the uncorrectable memory error because only the affected user jobs were terminated, and only the faulty page was offlined. Overall, considering both uncorrectable memory error recovery paths (RRE and error containment after an RRF), the impact of uncorrectable memory errors was alleviated 92% of the time on H100, while the GPU can remained operable. Such uninterrupted operations were not achievable on previous-generation GPUs (e.g., Kepler in (di2014lessons; tiwari2015reliability) and Volta in (oles2024understanding)), as an uncorrectable memory error would immediately cause user job interruption and put GPU in an error state, necessitating a GPU reset to recover (nv_xid_doc).
@@
Figure 6. Intra-GPU hardware error propagation probabilities in A100 GPUs. Numbers on the edges show propagation probabilities and average propagation time in seconds. :
@@
GSP errors can be caused by either GSP firmware bugs (nv_driver_release_notes) or demanding workload. For example, Delta SREs observed that these errors were highly correlated with heavy ML benchmarks, and they suggested that GSP errors are high-impact errors whose recovery requires manual node draining and reboots. Our propagation analysis confirm that the GSP is a single point of failure on both A100 GPUs in part because of their spontaneous nature and high downstream impact (e.g., GPU hangs) on the GPU.
@@
GPU Fallen Off the Bus. GPU Fallen Off the Bus errors were logged when the GPU driver could not reach the GPU over the system bus. This error is an integration error often caused by a loose GPU-motherboard connection or contact failure because of thermal cycles (tiwari2015reliability). Over 99% of the errors of this type lead to similar errors in close successions and eventually put the GPU into an error state.
Our GPU hardware error propagation analysis suggests that the error detection and recovery of GSP, PMU, and the communication interfaces (e.g., PMU SPI) need to be improved via duplications and error-detection and correction mechanisms to prevent single points of failure, as evident in A100 GPUs. In fact, AWS recommends disabling GSP for stability over performance benefits (aws_trouble). By improving GSP hardware and driver software combined with a tightly integrating CPU-GPU architecture, the H100 GPUs in the GH200 Superchips significantly improve GPU hardware resilience over the A100 GPUs. Notably, apart from three GSP errors, which follow the same propagation paths as GSP errors in A100 GPUs, we observed no other hardware error propagation paths on H100 GPUs.
@@
Figure 7. NVLink intra-GPU and inter-GPU error propagation in A100 GPUs. Numbers on the edges show propagation probabilities and average propagation time in seconds. :
@@
Figure 8. Incidents in which GPU errors led to job failure.: